OpenStudy (anonymous):

Find the sum of the following infinite geometric series if it exists 1/2 + -1/4 + 1/8 + -1/16

OpenStudy (anonymous):

We have x/y^2 . OK ?!

OpenStudy (anonymous):

But in each 2th we have (-) too !

OpenStudy (anonymous):

I know the common ratio is -1/2

OpenStudy (anonymous):

So what do you want ?!!!

OpenStudy (anonymous):

the sum of it

OpenStudy (anonymous):

In an arithmetic series I can find the sum of that but is there a formula for finding the sum of a geometric series?

OpenStudy (anonymous):

-0/1

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

YES ! LOOK :

OpenStudy (anonymous):

?....

OpenStudy (anonymous):

Number.(end.first)/2

OpenStudy (anonymous):

how do you find the last number?

OpenStudy (anonymous):

You have it ! You want a way to find the sum .HA? It s a way to find that .!

OpenStudy (anonymous):

There is no last number if it's infinite?

OpenStudy (anonymous):

I thought I was suppose to use sigma somehow?

OpenStudy (anonymous):

You want nth ? OK ?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

I just want to find the sum of the infinite geometric series if it's possible? -.-

OpenStudy (anonymous):

and I don't know how to

OpenStudy (anonymous):

I don't understand what you are trying do..

OpenStudy (anonymous):

In this you want to find the n th Number . Right ?

OpenStudy (anonymous):

what is the nth number?

OpenStudy (anonymous):

Amm for example when we say 2th day ... It s n th Number !

OpenStudy (anonymous):

No i dont want the nth number i want the sum of the entire infinite series..

OpenStudy (anonymous):

I can get the nth number but it asks for no nth number.

OpenStudy (anonymous):

it asks for the entire sum of the infinite geometric series.

OpenStudy (anonymous):

for a GP with first term 'a' and common ratio r less than 1 the sum for infinite terms is\[\frac{ a }{ 1-r } = S_n\] it is derived from applying the infinite limit to the formula for the sum of n terms.

OpenStudy (anonymous):

how do you get the sum of infinite terms?

OpenStudy (anonymous):

Does that equation equal the sum?..

OpenStudy (anonymous):

yes S_n is the sum of the infinite GP

OpenStudy (anonymous):

Then would it be 1/3?..

OpenStudy (anonymous):

Correct, the sum is \[\frac{ 1/2 }{ 1-(-1/2) } = \frac{1/2}{3/2}= \frac{1}{3}\]

OpenStudy (anonymous):

Okay. Thanks.